Could you please tell me how I would solve this question?:The sum of the squares of two consecutive positive odd integers is 202. find the integers.
Answer by stanbon(75887)

The sum of the squares of two consecutive positive odd integers is 202. find the integers.
----
1st odd: 2x-1
2nd odd: 2x+1
----
Equation:
(2x-1)^2 + (2x+1) = 202
-----
4x^2-4x+1 + 4x^2+4x+1 = 202
-----
8x^2+2 = 202
---
4x^2+1 = 101
4x^2 = 100
x^2 = 25
---
x = 5 or x = -5
---
If x = 5,
2x-1 = 9
2x+1 = 11
----
If x = -5
2x-1 = -11
2x+1 = -9
---------------------
Cheers,
Stan H.
================